OpenStudy (anonymous):

write the slope-intercept equation for the line pass through (-12, 10) and is perpendicular to 4x+6y=3.show work.

OpenStudy (dumbcow):

\[4x+6y =3 \rightarrow y = -\frac{2}{3}x + \frac{1}{2}\] perpendicular slope is opposite reciprocal, m = 3/2 use point to find y-intercept \[10 = \frac{3}{2}(-12) + b\] \[b=28\] \[y = \frac{3}{2}x + 28\]